Element X has two isotopes. If 92.22% of the element has an isotopic mass of 27.98 atomic mass units, and 7.78% of the element has an isotopic mass of 29.97 atomic mass units, the average atomic mass of element X is numerically equal to (rounded to two decimal places)


1
Expert's answer
2020-11-04T14:24:09-0500

Solution:

The average atomic mass of an element is the sum of the masses of its isotopes, each multiplied by its natural abundance.

Average atomic mass = f1M1 + f2M2 +… + fnMnwhere f is the fraction representing the natural abundance of the isotope and M is the mass number (weight) of the isotope.

Hence,

Average atomic mass of X = f1M1 + f2M2 = (0.9222 × 27.98 amu) + (0.0778 × 29.97 amu) = 28.1348 amu

Average atomic mass of X is 28.1348 = 28.13


Answer: the average atomic mass of element X is 28.13 amu.